Suppose in its income statement for the year ended June 30, 2022, The Clorox Company reported the following condensed data (dollars in millions).

Salaries and wages expenses

$ 460

Research and development expense

$ 114

Depreciation expense

90

Income tax expense

430

Sales revenue

5,830

Loss on disposal of plant assets

46

Interest expense

161

Cost of goods sold

2,800

Advertising expense

499

Rent expense

105

Sales returns and allowances

230

Utilities expense

60

(a)

Prepare a multiple-step income statement.

(b)

The parts of this question must be completed in order. This part will be available when you complete the part above.

(c1)

The parts of this question must be completed in order. This part will be available when you complete the part above.

(c2)

The parts of this question must be completed in order. This part will be available when you complete the part above.
Answer #1

The Multi-step income statement is shown below:

$460 $499 А Bc Multi-step income statement Particulars Amount Amout 3 Sales Revenue $5,830 4 Less: Sales return and allowance

The following excel formulas are used:

A 2 Amout =B3-B4 =C5-B6 Multi-step income statement Particulars Amount 3 Sales Revenue 5830 4 Less: Sales return and allowanc

Note:

The answer of part (a) is provided above and the other parts of the question are not provided in the question. Please make available the other parts (b) (c1) and (c2) also, so that I can provide the solution.